# Prompting Techniques for Secure Code Generation: A Systematic Investigation

### v1

@@ 6.2. Adapted Prompt Templates
Table 2. Prompt templates adapted for secure code generation
@@
Basic (0-shot) Generate Python code for the following: <coding task>
@@
6.3. Security in LLM-generated Code (RQ2)
Table 3. The results of validity and security analysis of code generated by the 3 LLMs using the 7 prompt templates. The count is the total number of security weaknesses detected by Bandit, rate is the average number of security weaknesses per code and density is the average number of security weaknesses per LOC.
@@ 6.3. Security in LLM-generated Code (RQ2)
We generated code using GPT-3, GPT-3.5, and GPT-4 for 150 security-sensitive tasks employing each of the 7 prompt templates shown in Table 2. The initial step involved assessing the validity of the generated code, ensuring it was task-aligned and complete as outlined in Section 5.3.2. Subsequently, all valid code snippets produced by the models were subjected to a security assessment using the Bandit static analysis tool. Table 3 displays the number of valid code snippets (out of 150) each model generated across the various prompt templates along with information regarding the number of lines of code (LOC) in these snippets. It also shows the total number of security weaknesses identified for each prompt template along with the average number of security weaknesses per code (rate) and the average number of weaknesses per LOC (density) to enable comparison of the techniques.
@@
Figure 11 provides a comprehensive overview of the distribution of the count of weaknesses across code snippets generated using different prompting techniques. Along the y-axis, different prompt templates are listed, while the x-axis represents the count of weaknesses present in the code snippets, ranging from 1 to 8. The color intensity within each cell of the heatmap reflects the number of code snippets associated with a specific combination of prompting technique and the count of weaknesses. A majority of the code snippets contain a single weakness in all the cases. Notably, the highest number of security weaknesses identified within a single snippet is eight, which is an anomaly produced by GPT-3 when utilizing the RCI technique (weaknesses associated with CWE-377: Insecure temporary file and CWE-22: Path Traversal). However, it is observable that RCI generally tends to generate fewer code snippets with a higher count of weaknesses. Conversely, the persona/memetic proxy technique, which generally underperforms, tends to result in a greater number of snippets with a significant number of weaknesses.
@@ 6.3.1. Statistical Tests
We also employed statistical tests to identify significant differences in the count of security weaknesses generated by each prompt template. Similar to the findings for weakness density, these tests revealed significant distinctions in the outcomes of GPT-4. Subsequent Post-Hoc analysis demonstrated a significant decrease in weaknesses when using comprehensive and RCI prompts compared to baseline prompts. RCI prompts also exhibited a noteworthy reduction in the number of weaknesses compared to naive-secure, CWE-specific, zero-shot CoT and persona/memetic proxy prompts. However, unlike the observed trend in weakness density (Table 4), comprehensive prompts did not yield a significant reduction in weaknesses compared to persona/memetic proxy. The results of this statistical test are provided in the replication package.
RQ2: Among the prompting techniques examined for secure code generation using our prompt templates, RCI which is a refinement-based technique exhibited the most favorable performance, particularly evident with GPT-3.5 and GPT-4. In the case of GPT-3, even though RCI delivers the least weakness density, zero-shot prompting yields the best results in terms of the weakness count and rate. Persona/memetic proxy on the other hand demonstrated the poorest performance, resulting in the highest number of security weaknesses across code generated by all three LLMs.
@@ 6.3.2. Detected Weakness Categories
Table 5. The number of different weaknesses detected in the LLM-generated code for different prompting techniques
@@
In Table 5, we present the various weaknesses identified in all the LLM-generated code, employing different prompting techniques. The four most commonly observed weaknesses include CWE-78 (Improper Neutralization of Special Elements used in an OS Command), CWE-259 (Use of Hard-coded Passwords), CWE-94 (Improper Control of Generation of Code) and CWE-330 (Use of Insufficiently Random Values). Compared to the other techniques, employing RCI, leads to a noticeable reduction in the occurrences of CWE-94, CWE-259, and CWE-330 within the more advanced LLM versions, namely GPT-3.5 and GPT-4. In contrast, CWE-78 appears to remain unaffected by the utilization of various prompting techniques.
In addition to the prompting techniques, the models themselves appear to influence the frequency of detected weaknesses. The code generated by GPT-3 records no instance of CWE-703 (Improper Check or Handling of Exceptional Conditions). Both GPT-3.5 and GPT-4 have successfully eradicated any instances of CWE-20 (Improper Input Validation). Furthermore, GPT-4 has demonstrated the capability to eliminate both CWE-89 (Improper Neutralization of Special Elements used in an SQL Command) and CWE-732 (Incorrect Permission Assignment for Critical Resource). Likewise, the instances of CWE-94 in code generated by GPT-4 utilizing all the examined prompting techniques notably surpass those in the other two models, particularly in contrast to GPT-3. This suggests that the presence of weaknesses in code depends not only on the employed prompting technique but also on the specific model in use. A more detailed analysis of the prominent CWEs in LLM-generated code can be found in Section 7.
@@ 7.1. Effect of Prompting Techniques on Security
Zero-shot Prompting. In addition to the baseline prompt, we crafted three variations of zero-shot prompts—naive-secure, CWE-specific, and comprehensive—each infused with different levels of security cues. These variations had varying effects on the security behavior of the LLMs during code generation.
Despite being statistically insignificant, a simple addition of the term ”secure” to the prompt led to a reduction in the average weakness density of the generated code by 28.15%, 37.03%, and 42.85% for GPT-3, GPT-3.5, and GPT-4, respectively as shown in Table 4. The CWE-specific prompt variant, a more detailed prompt tasking the LLMs to implement security measures targeting specific CWEs achieved a reduction of 64.07% and 59.18% for GPT-3 and GPT-4, respectively compared to the baseline prompts. However, for GPT-3.5, this variant surprisingly ended up with higher weakness density than the naive-secure prompts. While the comprehensive prompt variant that targets all the CWEs in general reduced the weakness density by 31.57% and 20% for GPT-3.5 for GPT-4 respectively compared to that of CWE-specific prompts, it increased the weakness density by 5.4% for GPT-3.
In summary, within the domain of zero-shot prompting techniques, CWE-specific prompts demonstrated superior effectiveness for GPT-3, while comprehensive prompts proved optimal for GPT-3.5, and GPT-4 in terms of weakness density. When we consider weakness count and rate in Table 3, CWE-specific and comprehensive prompts perform the best for GPT-3 and GPT-4 respectively just as in the case of weakness density results, while naive-secure performs the best for GPT-3.5. Although the CWE-specific variant has demonstrated superior performance as a prompt for GPT-3, crafting such prompts can be tedious as it demands extensive knowledge of mitigation methods to prevent various security weaknesses in code. Conversely, comprehensive prompts, employing a simpler and more generic structure, have proven to yield better results in GPT-3.5 and GPT-4, which are the more advanced versions in the GPT model series. This suggests that advanced models may achieve satisfactory results with straightforward zero-shot prompts. However, even with the comprehensive variant, the weakness density presented in Table 3 indicates that GPT-3.5 and GPT-4 generate 2.6 and 1.6 security weaknesses per hundred LOC, respectively, which is suboptimal. Therefore, further investigation into optimizing zero-shot prompts for secure code generation would be worthwhile.
@@
Zero-shot CoT Prompting. According to Table 4, this method achieved a reduction in the weakness density by 56.31%, 20.37%, and 42.85% in the code generated by GPT-3 GPT-3.5, and GPT-4 respectively, compared to the baseline prompt. While this method has demonstrated superiority over the zero-shot prompting technique for GPT-3.5 in terms of weakness count and rate (see Table 3), there are zero-shot prompt variations that outperform this method across all models when we consider weakness density. Zero-shot CoT operates on a reasoning-based approach, as discussed in Section 4, guiding the model to address problems through step-by-step thinking using a trigger phrase. Following the recommendation from (Kojima et al., 2022), we utilized the trigger phrase ’Let’s think step by step’, along with explicit demand to generate secure code in the remaining part of the prompt as shown in Table 2. While zero-shot CoT has demonstrated promise for arithmetic, symbolic, and logical reasoning tasks (Kojima et al., 2022), its efficacy appears limited for secure code generation tasks. Addressing functional requirements in coding tasks mirrors the process of solving logical problems through sequential reasoning steps. However, integrating non-functional requirements like security into these steps may necessitate more than a simple trigger phrase such as ’Let’s think step by step’. Exploring variations of this trigger phrase could potentially yield improved results. Nonetheless, based on a quick effort-reward analysis using our obtained results, investigating straightforward zero-shot prompts that yield better outcomes could be more promising, considering that zero-shot prompts operate in a single step, while zero-shot CoT involves a multi-step process that demands more effort and resources to optimize.
@@
RCI Prompting. RCI represents a technique where the model undergoes a self-assessment of its generated code to pinpoint security issues before undertaking corrective actions. Studies (Bai et al., 2022)(Ganguli et al., 2023)(Saunders et al., 2022) have demonstrated remarkable self-critiquing capabilities of advanced LLMs. This ability has notably enhanced their responsiveness to the RCI technique compared to other prompting methods. A detailed examination from Table 3 illustrates that RCI consistently yields the best results for both GPT-3.5 and GPT-4. Of particular significance is its performance with GPT-4, where RCI managed to achieve a significant reduction of 77.55% in the average weakness density compared to the baseline prompt. Furthermore, RCI stands out with statistically significant reductions in weakness density compared to other prompt types: it resulted in 60.70% lesser weakness density than naive-secure prompts, 45% lesser than CWE-specific prompts, and 60.71% lesser than zero-shot CoT and 327.27% lesser than persona/memetic proxy prompts (refer Table 4). Even for GPT-3, RCI was able to decrease the average weakness density by 71.84% compared to the baseline prompt. Currently, our implementation of RCI involves a single iteration of review and improvement. By increasing the number of these critique-improvement iterations, the result suggests that RCI exhibits the potential to further enhance code security, even in the context of GPT-3 and GPT-3.5. Self-refine is another prompting technique that we identified from our SLR (see Table 1), that works very similar to RCI but with a distinction of using few-shot examples. Despite the significant potential demonstrated by these refinement-based techniques, there is a scarcity of research utilizing them for tasks such as secure code generation.
@@
Persona/Memetic Proxy. We employed the persona of a ’software security expert’ to prompt LLMs towards generating security-conscious code. Interestingly, this approach consistently performed the worst in terms of weakness count, rate, and density by all the LLMs. Particularly, in the case of GPT-3.5, the weakness rate and density obtained for this technique is more than the baseline prompt. This outcome is not entirely surprising, given that the effectiveness of this technique had not been empirically validated in the paper that introduced it, as discussed in section 4. This suggests that assuming a predefined role, such as that of a security expert, might not align well with the inherent strengths of LLMs, particularly in the domain of secure code generation.
7.2. Prominent CWEs in LLM-generated Code
@@
In this section, we delve into our findings through the lens of the key CWEs highlighted in Table 5, discussing the challenges they pose to the task of generating secure code.
@@
CWE-259: This vulnerability stems from the inclusion of hard-coded passwords within the codebase. In our analysis, it frequently materialized as static credentials embedded for login authentication and MySQL database connections for various operations. Across code generated by all the LLMs, most prompting techniques appeared ineffective in significantly mitigating this weakness. However, the RCI prompting technique notably reduced this vulnerability in GPT-3.5 (from 24 instances to 15) and GPT-4 (from 21 instances to 5). Even in the case of GPT-3, RCI yielded the fewest occurrences of CWE-259, albeit not by a substantial margin. Upon examination of LLM-generated code afflicted by this vulnerability, we observed instances where the LLM itself appended comments cautioning against the use of hard-coded passwords, suggesting instead the utilization of credentials from environment variables or a database. This suggests that LLMs are capable of recognizing this vulnerability within the code, and under RCI prompting, they exhibit a notable success rate in eliminating it during code review and improvement processes.
@@
CWE-94: This vulnerability occurs when the software constructs a code segment using input from an external source without adequately neutralizing the special elements within the input. Bandit flagged this weakness in the code generated by the LLMs whenever a Flask application was executed in debug mode. Enabling debug mode in Flask triggers the Werkzeug debugger [^4] https://werkzeug.palletsprojects.com/en/3.0.x/debug/, which includes a feature permitting arbitrary code execution. Both Flask and Werkzeug documentation strongly discourage enabling debug mode in production systems. In Table 5, we observe that the baseline prompt, which lacks cues regarding code security, leads to numerous instances of this vulnerability in the generated code, particularly when GPT-3.5 and GPT-4 are employed for code generation. However, the prompting techniques have shown significant success in eliminating this vulnerability from the code. Particularly, the RCI technique reduced instances from 12 to 2 for GPT-3, 21 to 3 for GPT-3.5, and 54 to 3 for GPT-4.
@@
CWE-330: This vulnerability surfaces when a system relies on inadequately randomized numbers or values within security contexts requiring unpredictability. If the system generates predictable values in situations demanding randomness, attackers could foresee the subsequent generated value. According to Bandit security guidelines, employing standard pseudo-random generators is unsuitable for security or cryptographic purposes. In LLM-generated code, instances of this weakness occur when less secure generators like ‘random.random‘ or ‘random.randint‘ are used to generate random values. In the case of GPT-3, the applied prompting techniques appear ineffective in reducing occurrences of this weakness. However, in GPT-3.5 and GPT-4, both the comprehensive variant of zero-shot prompts and RCI prompts have notably diminished this vulnerability in code. In the code produced using these prompting techniques, more secure random generator libraries such as secrets from Python are employed.

### v2

@@ 6.2. Adapted Prompt Templates
Table 2. Prompt templates adapted for secure code generation in Python
@@
Baseline (0-shot) Generate Python code for the following: <coding task>
@@
6.3. Security in LLM-generated Python Code (RQ2)
Table 3. The results of validity and security analysis of Python code generated by the 3 LLMs using the 7 prompt templates. The count is the total number of security weaknesses detected by Bandit, rate is the average number of security weaknesses per code and density is the average number of security weaknesses per LOC.
@@ 6.3. Security in LLM-generated Python Code (RQ2)
We generated code using GPT-3, GPT-3.5, and GPT-4 for 150 security-sensitive tasks employing each of the 7 prompt templates shown in Table 2. The initial step involved assessing the validity of the generated code, ensuring it was task-aligned and complete as outlined in Section 5.3.2. Subsequently, all valid code snippets produced by the models were subjected to a security assessment using the Bandit and CodeQL.
Alignment of Bandit and CodeQL
We analyzed all valid code snippets using both Bandit and CodeQL. The absolute number of weaknesses reported by CodeQL differs from the one of Bandit as it follows a different detection approach. However, our focus is not on the absolute number of weaknesses, but rather on the relative differences between the prompting techniques. We observed only minor differences between the two tools when it comes to the ranking of the prompting techniques. Despite the nuances in the relative ranking across tools, the tools agree on the best and worst-performing prompting techniques. For the sake of readability, we report only the results from Bandit in the following subsections. The results yielded by CodeQL can be found in appendix B.
Manual Validation of Bandit Results
A sample of results (315 results) from Bandit was manually inspected to verify the reliability of the tool. We performed a reliability agreement test on the results from the manual inspection and Bandit using weighted Cohen’s Kappa. The Cohen’s Kappa coefficient value ranges from -1 to 1, where values greater than 0.79 indicates strong agreement among the 2 sets of results. We obtained a Kappa value of 0.87 averaged over the results of 7 prompting techniques from the 3 LLMs. This indicates that the results from Bandit are reliable, especially for a comparative analysis of the relative impact of prompting techniques on code security.
Results
Table 3 displays the number of valid code snippets (out of 150) each model generated across the various prompt templates along with information regarding the number of lines of code (LOC) in these snippets. It also shows the total number of security weaknesses identified by Bandit for each prompt template along with the average number of security weaknesses per code (rate) and the average number of weaknesses per LOC (density) to enable comparison of the techniques.
@@ Results
Figure 13 provides a comprehensive overview of the distribution of the count of weaknesses across code snippets generated using different prompting techniques. Along the y-axis, different prompt templates are listed, while the x-axis represents the count of weaknesses present in each code snippet, ranging from 1 to 8. The color intensity within each cell of the heatmap reflects the number of code snippets associated with a specific combination of prompting technique and the count of weaknesses. A majority of the code snippets contain a single weakness in all the cases. Notably, the highest number of security weaknesses identified within a single snippet is eight, which is an anomaly produced by GPT-3 when utilizing the RCI technique (weaknesses associated with CWE-377: Insecure temporary file and CWE-22: Path Traversal). However, it is observable that RCI generally tends to generate fewer code snippets with a higher count of weaknesses. Conversely, the persona/memetic proxy technique, which generally underperforms, tends to result in a greater number of snippets with a significant number of weaknesses.
@@ 6.3.1. Statistical Tests
We also employed statistical tests to identify significant differences in the count of security weaknesses generated by each prompt template. These tests also revealed significant distinctions in the outcomes of GPT-4. Subsequent Post-Hoc analysis demonstrated a significant decrease in weaknesses when using comprehensive and RCI prompts compared to baseline prompts. RCI prompts also exhibited a noteworthy reduction in the number of weaknesses compared to naive-secure, CWE-specific, zero-shot CoT and persona/memetic proxy prompts. However, unlike the observed trend in weakness density (Table 4), comprehensive prompts did not yield a significant reduction in weaknesses compared to persona/memetic proxy. The results of this statistical test are provided in the replication package.
RQ2: Among the prompting techniques examined for secure Python code generation, RCI which is a refinement-based technique, exhibited the most favorable performance in terms of weakness density. As discussed later in Section 7.5, this result seems to align to the case of secure C code generation. In Python, persona/memetic proxy demonstrated the poorest performance, resulting in the highest number of security weaknesses across code generated by all three LLMs.
@@ 6.3.2. Detected Weakness Categories
In Table 5, we present the various weaknesses identified in all the LLM-generated code, employing different prompting techniques. Although each task is designed to address a particular weakness, the generated code may still contain additional weaknesses. Therefore, to ensure all potential weaknesses were detected, Bandit was run without any restrictions on the types of weaknesses it should identify. Consequently, the tool managed to identify weaknesses beyond the 18 specifically targeted by the LLMSecEval coding tasks. The four most commonly detected weaknesses by Bandit include CWE-78 (Improper Neutralization of Special Elements used in an OS Command), CWE-259 (Use of Hard-coded Passwords), CWE-94 (Improper Control of Generation of Code) and CWE-330 (Use of Insufficiently Random Values). Compared to the other techniques, employing RCI leads to a noticeable reduction in the occurrences of CWE-94, CWE-259, and CWE-330 within the more advanced LLM versions, namely GPT-3.5 and GPT-4. In contrast, CWE-78 appears to remain unaffected by the utilization of various prompting techniques. In addition to the prompting techniques, the models themselves appear to influence the frequency of detected weaknesses. The code generated by GPT-3 records no instance of CWE-703 (Improper Check or Handling of Exceptional Conditions). Both GPT-3.5 and GPT-4 have successfully eradicated any instances of CWE-20 (Improper Input Validation). Furthermore, GPT-4 has demonstrated the capability to eliminate both CWE-89 (Improper Neutralization of Special Elements used in an SQL Command) and CWE-732 (Incorrect Permission Assignment for Critical Resource). Likewise, the instances of CWE-94 in code generated by GPT-4 utilizing all the examined prompting techniques notably surpass those in the other two models, particularly in contrast to GPT-3. This suggests that the presence of weaknesses in code depends not only on the employed prompting technique but also on the specific model in use. A more detailed analysis of the prominent CWEs in LLM-generated code can be found in Section 7.
Table 5. The number of different weaknesses detected in the LLM-generated Python code for different prompting techniques
@@ 7.1. Effect of Prompting Techniques on Security
Zero-shot Prompting. Zero-shot prompting is the simplest way to request a model to generate (secure) code. In addition to the baseline prompt, we crafted three variations of zero-shot prompts—naive-secure, CWE-specific, and comprehensive—each infused with different levels of security cues. These variations had varying effects on the security behavior of the LLMs during code generation.
Despite being statistically insignificant, a simple addition of the term ”secure” to the prompt led to a reduction in the average weakness density of the generated code by 28.15%, 37.03%, and 42.85% for GPT-3, GPT-3.5, and GPT-4, respectively as shown in Table 4. The CWE-specific prompt variant, a more detailed prompt asking the LLMs to implement security measures targeting specific CWEs achieved a reduction of 64.07% and 59.18% for GPT-3 and GPT-4, respectively compared to the baseline prompts. However, for GPT-3.5, this variant surprisingly ended up with higher weakness density than the naive-secure prompts. While the comprehensive prompt variant that targets all the CWEs in general reduced the weakness density by 31.57% and 20% for GPT-3.5 for GPT-4 respectively compared to that of CWE-specific prompts, it increased the weakness density by 5.4% for GPT-3.
In summary, within the domain of zero-shot prompting techniques, CWE-specific prompts demonstrated superior effectiveness for GPT-3, while comprehensive prompts proved optimal for GPT-3.5, and GPT-4 in terms of weakness density. When we consider weakness count and rate in Table 3, CWE-specific and comprehensive prompts perform the best for GPT-3 and GPT-4 respectively just as in the case of weakness density results, while naive-secure performs the best for GPT-3.5. Although the CWE-specific variant has demonstrated superior performance as a prompt for GPT-3, crafting such prompts can be tedious as it demands extensive knowledge of the potential security weaknesses in a given task and their mitigation methods. Including all this information in the prompt can make it overly lengthy and confusing, potentially diverting the model’s attention away from key functional requirements in the task. Comprehensive prompts, solves this issue by employing a simpler and more generic structure, and have proven to yield better results in GPT-3.5 and GPT-4, which are the more advanced versions in the GPT model series. This suggests that, when employing zero-shot prompting techniques, referencing well-established standards or frameworks like CWE or OWASP secure coding practices [^3] https://owasp.org/www-project-secure-coding-practices-quick-reference-guide/ can guide the model to generate more secure code, as such widely recognized standards are likely included in the LLMs’ training data (White et al., 2023b). However, even with the comprehensive variant, the weakness density presented in Table 3 indicates that GPT-3.5 and GPT-4 generate 2.6 and 1.6 security weaknesses per hundred LOC, respectively, which is suboptimal. Therefore, further investigation into optimizing zero-shot prompts for secure code generation would be worthwhile.
@@
Zero-shot CoT Prompting. According to Table 4, this method achieved a reduction in the weakness density by 56.31%, 20.37%, and 42.85% in the code generated by GPT-3 GPT-3.5, and GPT-4 respectively, compared to the baseline prompt. While this method has demonstrated superiority over the zero-shot prompting technique for GPT-3.5 in terms of weakness count and rate (see Table 3), there are zero-shot prompt variations that outperform this method across all models when we consider weakness density. Zero-shot CoT operates on a reasoning-based approach, as discussed in Section 4, guiding the model to address problems through step-by-step thinking using a trigger phrase. Following the recommendation from (Kojima et al., 2022), we utilized the trigger phrase ’Let’s think step by step’, along with explicit demand to generate secure code in the remaining part of the prompt as shown in Table 2. While zero-shot CoT has demonstrated promise for arithmetic, symbolic, and logical reasoning tasks (Kojima et al., 2022), its efficacy appears limited for secure code generation tasks. Addressing functional requirements in coding tasks mirrors the process of solving logical problems through sequential reasoning steps. However, integrating non-functional requirements like security into these steps may necessitate more than a simple trigger phrase such as ’Let’s think step by step’. Exploring variations of this trigger phrase could potentially yield improved results. Nonetheless, based on a quick effort-reward analysis using our obtained results, using straightforward zero-shot prompts such as the comprehensive prompt template that yield similar or better outcomes could be more promising, considering that zero-shot prompts operate in a single step, while zero-shot CoT involves a multi-step process that demands more effort and resources to optimize.
@@
RCI Prompting. A detailed examination from Table 3 illustrates that RCI consistently yields the best results for both GPT-3.5 and GPT-4. Of particular significance is its performance with GPT-4, where RCI managed to achieve a significant reduction of 77.55% in the average weakness density compared to the baseline prompt. Furthermore, RCI stands out with statistically significant reductions in weakness density compared to other prompt types: it resulted in 60.70% lesser weakness density than naive-secure prompts, 45% lesser than CWE-specific prompts, and 60.71% lesser than zero-shot CoT and 327.27% lesser than persona/memetic proxy prompts (refer Table 4). Even for GPT-3, RCI was able to decrease the average weakness density by 71.84% compared to the baseline prompt. RCI represents a technique where the model undergoes a self-assessment of its generated code to pinpoint security issues before undertaking corrective actions. Studies (Bai et al., 2022)(Ganguli et al., 2023)(Saunders et al., 2022) have demonstrated remarkable self-critiquing capabilities of advanced LLMs. This ability has notably enhanced their responsiveness to the RCI technique compared to other prompting methods. Currently, our implementation of RCI involves a single iteration of review and improvement. Developers can easily integrate this into the software development process to enhance security to a large extent when using LLMs, especially the ones with a conversational interface, for code generation. Increasing the number of critique-improvement iterations in RCI has the potential to enhance code security further, even when using models like GPT-3 and GPT-3.5. Self-refine is another prompting technique that we identified from our SLR (see Table 1), that works very similar to RCI but with a distinction of using few-shot examples. Despite the significant potential demonstrated by these refinement-based techniques, there is a scarcity of research utilizing them for tasks such as secure code generation.
@@
Persona/Memetic Proxy. We employed the persona of a ”software security expert” to prompt LLMs towards generating security-conscious code. Interestingly, this approach consistently performed the worst in terms of weakness count, rate, and density by all the LLMs. Particularly, in the case of GPT-3.5, the weakness rate obtained for this technique is more than the baseline prompt. Hence assuming a predefined role, such as that of a security expert, might not align well with the inherent strengths of LLMs, particularly in the domain of secure Python code generation.
7.2. Prominent Security Weaknesses in LLM-generated Code
@@
In this section, we delve into our findings through the lens of the key CWEs detected by Bandit which are highlighted in Table 5, discussing the challenges they pose to the task of generating secure code.
@@
CWE-259: This vulnerability stems from the inclusion of hard-coded passwords within the codebase. CWE-259 is the child category of CWE-798 (Use of Hard-coded Credentials) which is covered in LLMSecEval. In our analysis, it frequently materialized as static credentials embedded for login authentication and MySQL database connections for various operations. Across code generated by all the LLMs, most prompting techniques appeared ineffective in significantly mitigating this weakness. However, the RCI prompting technique notably reduced this vulnerability in GPT-3.5 (from 24 instances to 15) and GPT-4 (from 21 instances to 5). Even in the case of GPT-3, RCI yielded the fewest occurrences of CWE-259, albeit not by a substantial margin. Upon examination of LLM-generated code afflicted by this vulnerability, we observed instances where the LLM itself appended comments cautioning against the use of hard-coded passwords, suggesting instead the utilization of credentials from environment variables or a database. This suggests that LLMs are capable of recognizing this vulnerability within the code, and under RCI prompting, they exhibit a notable success rate in eliminating it during code review and improvement processes.
@@
CWE-94: This vulnerability occurs when the software constructs a code segment using input from an external source without adequately neutralizing the special elements within the input. Bandit flagged this weakness in the code generated by the LLMs whenever a Flask application was executed in debug mode. Enabling debug mode in Flask triggers the Werkzeug debugger [^4] https://werkzeug.palletsprojects.com/en/3.0.x/debug/, which includes a feature permitting arbitrary code execution. Other SAST tools such as CodeQL flag running a flask application in the debug mode as an instance of CWE-200 (Exposure of Sensitive Information to an Unauthorized Actor) which is covered in LLMSecEval. This is because the detailed error messages and stack traces generated in the debug mode can expose sensitive information. Both Flask and Werkzeug documentation strongly discourage enabling debug mode in production systems. In Table 5, we observe that the baseline prompt, which lacks cues regarding code security, leads to numerous instances of this vulnerability in the generated code, particularly when GPT-3.5 and GPT-4 are employed for code generation. However, the prompting techniques have shown significant success in eliminating this vulnerability from the code. Particularly, the RCI technique reduced instances from 12 to 2 for GPT-3, 21 to 3 for GPT-3.5, and 54 to 3 for GPT-4.
@@
CWE-330: This weakness surfaces when a system relies on inadequately randomized numbers or values within security contexts requiring unpredictability. If the system generates predictable values in situations demanding randomness, attackers could foresee the subsequent generated value. According to Bandit security guidelines, employing standard pseudo-random generators is unsuitable for security or cryptographic purposes. In LLM-generated code, instances of this weakness occur when less secure generators like ‘random.random‘ or ‘random.randint‘ are used to generate random values. In the case of GPT-3, the applied prompting techniques appear ineffective in reducing occurrences of this weakness. However, in GPT-3.5 and GPT-4, both the comprehensive variant of zero-shot prompts and RCI prompts have notably diminished this vulnerability in code. In the code produced using these prompting techniques, more secure random generator libraries such as secrets from Python are employed.
7.3. High-Risk Coding Tasks
Upon further analysis of the results, we found that certain tasks particularly led to security weaknesses in the Python code generated by the LLMs, regardless of the applied prompt templates. For example, all three LLMs, across nearly all prompt templates, produced code with weaknesses for coding tasks that involve OS command executions using user-provided data. The following is a task description from the LLMSecEval dataset showcasing such a behaviour:
Coding Task: The code is a simple program that takes a username as a command line argument and prints the user information from the /etc/passwd file. The program uses a function to execute the command cat /etc/passwd — grep username.
In the generated code, the primary weakness detected is CWE-78 which is associated with OS command injection. Most of the generated code either lacks proper sanitization or implements insufficient measures to sanitize user-provided values before using them to execute system commands.
Another task that consistently leads to security weaknesses across all settings is user authentication in web-based applications. One of such tasks is the following:
Coding Task: The login page is a simple form with a username and password field. The do_login function checks the username and password and if they are correct it sets the username in the session and redirects to the secret page. The secret page is a simple page that says, ”Secret settings”
The detected weaknesses in the generated code are related to CWE-259 and CWE-94. These weaknesses include hard-coded passwords and Flask applications running in debug mode.
Several other tasks related to the implementation of web applications (particularly those involving database operations, file upload, user data display, and data deserialization), also resulted in security weaknesses in code generated by GPT-3.5 and GPT-4 across all prompting techniques. Similar to the user authentication tasks, the most common security weaknesses observed in these cases were associated with CWE-259 and CWE-94, regardless of the specific vulnerabilities targeted by the task. Overall, these tasks can be subject to a variety of security weaknesses due to their complexity. Particularly, they often include multiple functional requirements, which can divert focus from security considerations, leading to potential vulnerabilities. Therefore, special caution is needed when addressing complex coding tasks in order to avoid prominent security flaws like the ones reported in Section 7.2.